The length of a rectangle is 3 in. more than twice its width. If the perimeter of the rectangle is 36 in., find the width of the rectangle.
Answer by checkley71(8403)

L=2W+3
2(2W+3)+2W=36
4W+6+2W=36
6W=36-5
6W=30
W=30/6
W=5 in.
PROOF
2(2*5+3)+2*5=36
2(10+3)+10=36
2*13+10=36
26+10=36
36=36